OpenStudy (anonymous):

HHEEEEEELLLLLLLLPPPPPPPP. find the coordinates of the vertex and the x and y intercepts. y=x^2-3x+2

OpenStudy (anonymous):

differentiate wrt x and equate to 0 to get the maxima

OpenStudy (anonymous):

can you lead me step by step please? i know the answer is (vertex: 3/2, -1/4) and x-intercepts are (2,0) and (1,0) and y-intercepts are (0,2) but idk how to get those answers

OpenStudy (anonymous):

first differentiate equation to get y' =2x-3 equate it to 0 0=2x-3 x=3/2 now plug this x in original equation to get y

OpenStudy (anonymous):

wait where did you get y=2x-3?

OpenStudy (anonymous):

do u know differenciation, i differenciated the first equation

OpenStudy (anonymous):

ive never even heard of that what do you do?

OpenStudy (anonymous):

i'm sorry...............